OpenStudy (anonymous):

Which equation is a quadratic equation? A.) 3x − y = 8x + 2 B.)y = 7x(5 − 2x) + 9 C.) 7x2 + y = (3 + x)(7x − 1) D.) y − 3 = (x2 − 3)(8x + 3)

OpenStudy (neo92):

a quadratic eqn is a second degree polynomial with a single variable. which can be genarally expressed as, \[ax^{2}+bx+c=0\text{ where a,b,c are constants and }a\neq0\] can u do it now?

OpenStudy (neo92):

try and tel me ur answer

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

im confused?

OpenStudy (anonymous):

is it C..?

OpenStudy (neo92):

is y in this problem is a constant o not?

OpenStudy (neo92):

@zakiaboo if y is not a constant.. theres a little problem..

OpenStudy (anonymous):

umm i think so.

OpenStudy (amistre64):

C has other issues with it other than a y

OpenStudy (anonymous):

@zakiaboo try using what @Neo92 said ax^2+bx+c=0 if you put some of your answers in to that form it may help you figure it out.

OpenStudy (amistre64):

7x2 + y = (3 + x)(7x − 1) 7x2 + y = 7x^2 + (other stuff) since both sides contain 7x^2, its not really there ..... 7x2 + y = 7x^2 + (other stuff) -7x^2 -7x^2 ----------------------- y = (other stuff)

OpenStudy (neo92):

@amistre64 yep.. it got no terms of 2nd order

OpenStudy (anonymous):

oh you mean put it in the order that its suppose to be in..?

OpenStudy (amistre64):

C turns out to be a trick that they are playing on you; it LOOKS quadratic, but its really linear

OpenStudy (anonymous):

yea i just figured out its B.

OpenStudy (anonymous):

right?

OpenStudy (amistre64):

b is good

OpenStudy (amistre64):

y = 7x(5 − 2x) + 9 y = 14x^2 + .....

OpenStudy (anonymous):

yay for B